# SplatSuRe: Selective Super-Resolution for Multi-view Consistent 3D Gaussian Splatting

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) enables high-quality novel view synthesis, motivating interest in generating higher-resolution renders than those available during training. A natural strategy is to apply super-resolution (SR) to low-resolution (LR) input views, but independently enhancing each image introduces multi-view inconsistencies, leading to blurry renders. Prior methods attempt to mitigate these inconsistencies through learned neural components, temporally consistent video priors, or joint optimization on LR and SR views, but all uniformly apply SR across every image. In contrast, our key insight is that close-up LR views may contain high-frequency information for regions also captured in more distant views, and that we can use the camera pose relative to scene geometry to inform where to add SR content. Building from this insight, we propose SplatSuRe, a method that selectively applies SR content only in undersampled regions lacking high-frequency supervision, yielding sharper and more consistent results. Across Tanks & Temples, Deep Blending and Mip-NeRF 360, our approach surpasses baselines in both fidelity and perceptual quality. Notably, our gains are most significant in localized foreground regions where higher detail is desired.
Figure 1: SplatSuRe trains a 3D Gaussian Splatting [kerbl3Dgaussians] model to produce sharp, high-resolution novel views from low-resolution inputs. By selectively leveraging high-frequency cues already present in low-resolution training views and applying super-resolution only where needed, our method delivers greater detail and multi-view consistency than prior approaches without any additional training.
@@ 1 Introduction
Novel view synthesis aims to render unseen viewpoints given a set of multi-view images. 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] enables real-time, photorealistic novel view synthesis by representing scenes as an explicit set of anistropic Gaussians optimized through differentiable splatting. While 3DGS excels at efficiency and reconstruction quality, its performance is tightly coupled to the resolution of the training images. Models trained on low-resolution (LR) inputs lack access to high-frequency signals present in high-resolution (HR) views, resulting in blurry textures, over-smoothed surfaces, and aliasing artifacts when rendered at higher test-time resolutions [Yu2024MipSplatting].
When restricted to only LR views, a natural strategy is to apply super-resolution (SR) to enhance them before fitting a 3D model [feng2024srgssuperresolution3dgaussian]. However, single-image SR operates independently on each view and frequently introduces view-dependent hallucinated textures [wang2024exploiting]. When these inconsistent SR predictions are used as direct supervision, the resulting 3D optimization receives conflicting gradients across viewpoints, degrading model quality. Prior methods attempt to mitigate these inconsistencies through learned neural components, temporally consistent video priors, or joint optimization on LR and SR views [Shen2024SuperGaussian, wan2025s2gaussian, xie2024supergssuperresolution3dgaussian, huang2024assrnerfarbitraryscalesuperresolutionvoxel, yu2024gaussiansr]. However, these methods inject SR content uniformly across the image, regardless of whether a region actually benefits from generative detail or is already well-constrained by existing LR observations.
In contrast, the key observation motivating our work is that images of a scene do not sample 3D content uniformly. A low-resolution view captured up close often contains enough high-frequency detail to supervise rendering of more distant views that observe the same region only coarsely. This disparity in multi-view sampling implies that many views already receive high-frequency supervision from closer views, whereas others that do not have any closer views providing sufficient higher-frequency information would benefit from SR guidance. Applying SR indiscriminately therefore introduces unnecessary inconsistencies in well-resolved regions.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] enables real-time novel view synthesis by representing scenes as sets of anisotropic Gaussians optimized through differentiable rasterization. While it achieves high rendering efficiency, models trained with low resolution (LR) images suffer from aliasing artifacts when rendered at higher resolutions, since Gaussians optimized at coarse scales become undersampled. Mip-Splatting [Yu2024MipSplatting] mitigates this aliasing by applying scale-adaptive 3D and 2D filtering while preserving radiance energy across resolutions, similar in spirit to Mip-NeRF [barron2022mipnerf360]. While achieving significant improvement over vanilla 3DGS, its rendering quality is still tied to the resolution of training views and blurring can occur at higher resolutions. In contrast, our method uses super-resolution to further inform high-resolution information.
@@
Super-resolution (SR) has long been applied to neural radiance fields [mildenhall2021nerf] to enhance novel view synthesis quality [wang2022nerf, 10205402, feng2023zssrtefficientzeroshotsuperresolution, vishen2025advancingsuperresolutionneuralradiance, zheng2025supernerfganuniversal3dconsistentsuperresolution, huang2024assrnerfarbitraryscalesuperresolutionvoxel, roessle2023ganerf]. More recently, SR has been extended to 3D Gaussian Splatting (3DGS). Several variants aim to recover high-resolution detail and improve multi-view consistency [feng2024srgssuperresolution3dgaussian, Shen2024SuperGaussian, wan2025s2gaussian, xie2024supergssuperresolution3dgaussian] through residual feature learning [xie2024supergssuperresolution3dgaussian], uncertainty modeling [wan2025s2gaussian, xie2024supergssuperresolution3dgaussian], per-scene refinement [wan2025s2gaussian] or video super-resolution [Shen2024SuperGaussian]. Among these, SRGS [feng2024srgssuperresolution3dgaussian] jointly optimizes Gaussian parameters using both LR ground-truth images and super-resolved views produced by a frozen single-image SR model. However, applying this enhancement uniformly across all regions does not eliminate the effect of inconsistencies introduced by super-resolution. S2Gaussian [wan2025s2gaussian] focuses on sparse view reconstruction and proposes an inconsistency modeling module trained per-scene to reduce inconsistencies in SR images. SuperGaussian [Shen2024SuperGaussian] applies a video SR network to frames rendered from a low-resolution 3DGS model, using the resulting temporally consistent sequence for retraining the 3D model. While these approaches improve fidelity, they either rely on additional neural components to enforce consistency or lack spatial adaptivity in the use of SR. In contrast, our method utilizes camera pose information relative to scene geometry to determine undersampled regions, and selectively applies super-resolution in those regions to enhance sharpness while maintaining fidelity.
@@
Diffusion methods offer complementary advances in 3D representation and enhancement tasks. Several works jointly optimize diffusion and 3D parameters for view-consistent generation and super-resolution [lin2025diffsplat, DiffGS, yu2024gaussiansr, chen2025bridgingdiffusionmodels3d]. GaussianSR [yu2024gaussiansr] distills information from 2D super-resolution models as a loss for training a 3DGS model. 3DSR [chen2025bridgingdiffusionmodels3d] uses a 3DGS model to enforce 3D consistency in the diffusion process, fitting it multiple times through denoising diffusion steps. Diffusion priors have also been applied as post-processing to enhance renders from 3DGS models [liu20243dgs, wu2025difix3d, gsfix3d]. While these methods rely on large pretrained diffusion models to predict images that reduce 3D inconsistencies, we leverage explicit geometric relationships between cameras and scene structure to determine where generative detail is needed, which can then be added via any of these complementary methods.
@@
In addition to image and video-based models, SR can also be performed directly in 3D, circumventing multi-view inconsistencies. Geometric point-based networks [yu2018pu, Liu_2024_CVPR, fang2024egp3dedgeguidedgeometricpreserving, dinesh20193dpointcloudsuperresolution] locally upscale point clouds to recover fine-grained geometry. However, these approaches typically only upsample geometry, while texture resolution still relies on 2D methods. As before, our method is orthogonal to these techniques and can be combined with them to further improve reconstruction fidelity.
@@ 3 Background
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] models a scene as a set of 3D Gaussians, each represented by a mean position $\boldsymbol{\mu}\in\mathcal{R}^{3}$, per-axis scale $\boldsymbol{s}\in\mathcal{R}^{3}$, rotation vector $\boldsymbol{q}\in\mathcal{R}^{4}$, scalar opacity $o\in\mathcal{R}_{+}$, and view-dependent color $\boldsymbol{c}$, represented as a base color with spherical harmonic coefficients. Volumetric rendering with alpha blending is used to splat the Gaussians onto the image plane and the resulting color for pixel $\boldsymbol{p}$ is given by:
@@
where $N$ is the number of Gaussians intersecting the pixel ray, $\alpha^{i}(\boldsymbol{p})=o^{i}e^{-\frac{1}{2}(\boldsymbol{p}-\boldsymbol{\mu}^{i})^{T}(\boldsymbol{\Sigma}_{{2D}}^{i})^{-1}(\boldsymbol{p}-\boldsymbol{\mu}^{i})}$ is the contribution of the $i^{th}$ Gaussian, and $\boldsymbol{c}^{i}$ is its view-dependent color. The 2D covariance matrix $\boldsymbol{\Sigma}_{2D}$ is given by the EWA Splatting approximation [zwicker2002ewa]. Gaussians are sorted in depth order before splatting to ensure that transmittance is computed correctly.
@@ 3.2 Super-Resolution for Gaussian Splatting
When trained solely on low-resolution (LR) images, 3DGS models lack access to the high-frequency cues present in high-resolution (HR) views, leading to over-smoothed textures and incomplete fine structure. This limitation has motivated the integration of super-resolution (SR) models into the 3DGS training pipeline. SRGS [feng2024srgssuperresolution3dgaussian] employs a frozen single-image SR model to generate super-resolved views and jointly optimizes Gaussian parameters against both LR ground truth and SR outputs, supplying explicit high-frequency supervision that LR views alone cannot provide. However, SRGS applies SR uniformly across the entire image, even in regions that already receive reliable high-frequency supervision from nearby LR views. Since SR is applied independently to each image, the generated details are not necessarily multi-view consistent – injecting them everywhere can introduce geometric or texture inconsistencies, leading to averaging effects in the model that render as blurring.
@@ 4.1 Gaussian Fidelity Score
A high ratio indicates that the Gaussian is sampled at varied frequencies across views, meaning some views observe it with high fidelity and can supervise the others. Conversely, a ratio close to one indicates uniformly sampling across views, indicating that this region lacks any higher-frequency observations and requires SR to add generated details. This interpretation also holds for view regions where the Gaussian projects near its maximal radius, since no other views provide higher frequency information.
@@ 4.2 Super-Resolution Region Selection
$\mathcal{M}(t)$ $=\{\mathcal{G}^{i}\ |\ t=\operatorname*{argmax}_{t\in T}r^{i}_{t}\;\forall i\in N\},$ (8)
@@
where Render()($\cdot$) denotes splatting the LR model by alpha-blending the specified per-Gaussian values, $\mathcal{G}$ is the set of all Gaussians, $\mathbf{score}_{\mathcal{G}}$ is the vector of Gaussian fidelity scores, and $\mathbf{1}_{\mathcal{M}(t)}(\mathcal{G})$ is an indicator that is 1 for Gaussians in $\mathcal{M}(t)$ and 0 otherwise. The first term in Equation 9 ensures that undersampled regions with low fidelity scores receive SR, while the second term ensures that areas observed most closely by the current view also receive SR because no alternative view provides higher resolution information. Finally, the weight map is normalized to ensure the magnitude of the SR loss is consistent across views. As illustrated in Figure 4, this per-view weight map is high in regions requiring SR and low in regions already sufficiently supervised by LR views.
@@ 4.3 SplatSuRe Training Objective
In each training iteration, the model is rendered at the target high resolution. The render $R_{HR}$ is downsampled to produce $R_{HR}\downarrow$ and compared with the LR ground truth $I_{LR}$, yielding the LR loss:
@@
Figure 5: Qualitative results on Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018] and Mip-Nerf 360 [barron2022mipnerf360]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [Yu2024MipSplatting] and SRGS [feng2024srgssuperresolution3dgaussian], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves finer details in text (red box on truck), high-frequency patterns (yellow box on carpet and green box on tray) and recognizable distant objects observed in other views (blue box on church mural). Notably, it reduces Gaussian artifacts (orange arrow) observed in other methods. Additional results in Appendix A.3.
@@ 5 Experiments
We integrate our SplatSuRe method into the 3D Gaussian Splatting (3DGS) codebase [liu20243dgs], adding modules for true Gaussian radius computation, weight map computation and rendering, and auxiliary loss terms. Experiments are conducted at $4\times$ super-resolution using StableSR [wang2024exploiting] and using a ratio threshold $\tau{=}1.1$ to generate weight maps. Our choice of ratio threshold and SR model are ablated in Sections 7.1 and 7.2, respectively. We use $\lambda{=}0.2$ and $\gamma{=}0.4$ for our losses defined in Equations 10-12.
@@
Table 1: Quantitative results on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. The best, second best and third best entries are highlighted. Our SplatSuRe method achieves the strongest results across all metrics.
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 0.669 19.41 0.350 71.58 2.013 0.0895 57.776 3.412
StableSR [wang2024exploiting] 0.751 22.47 0.300 59.29 1.123 0.0667 56.748 4.945
Mip-Splatting [Yu2024MipSplatting] 0.767 23.10 0.303 52.46 1.137 0.0597 46.571 5.043
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.771 23.32 0.286 49.11 1.048 0.0535 55.209 4.633
Ours + StableSR [wang2024exploiting] 0.784 23.81 0.272 37.72 1.040 0.0413 58.332 3.928
Table 2: Quantitative results on Deep Blending [DeepBlending2018] and Mip-NeRF 360 [barron2022mipnerf360]. Our SplatSuRe method achieves the strongest results across all metrics on Deep Blending and outperforms SRGS [feng2024srgssuperresolution3dgaussian] on Mip-NeRF 360. Appendix A.1 and A.2 present $8\times$ SR and per-scene results.
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 0.836 26.72 0.335 0.868 0.0574 0.635 20.67 0.384 0.645 0.0529
StableSR [wang2024exploiting] 0.845 27.16 0.325 0.630 0.0505 0.699 24.28 0.351 0.725 0.0387
Mip-Splatting [Yu2024MipSplatting] 0.865 28.43 0.327 0.690 0.0398 0.759 26.48 0.292 0.183 0.0132
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.861 28.23 0.317 0.630 0.0409 0.734 25.92 0.317 0.339 0.0194
Ours + StableSR [wang2024exploiting] 0.872 29.01 0.306 0.496 0.0330 0.740 26.34 0.323 0.339 0.0179
@@
We compare SplatSuRe with four representative baselines using their official implementations on our evaluation datasets. 3DGS (LR) [kerbl3Dgaussians] is trained on low-resolution inputs and serves as the primary baseline. StableSR [wang2024exploiting] applies a single-image SR model to the low-resolution training data and fits a 3DGS model on the super-resolved images. Mip-Splatting [Yu2024MipSplatting] is trained at low resolution and mitigates aliasing through 3D and 2D multi-scale filtering for high-resolution rendering. Finally, SRGS [feng2024srgssuperresolution3dgaussian] optimizes Gaussian parameters using both low-resolution and super-resolved images from a frozen SR model, without additional pretraining or scene-specific fine-tuning.
@@
We evaluate both reconstruction fidelity and perceptual realism using eight complementary metrics. For reference-based assessment, we report SSIM, PSNR, and LPIPS [zhang2018perceptual], which measure pixel and feature-level agreement with ground-truth high-resolution images. To capture distributional and semantic perceptual quality, we include FID [heusel2017gans], CMMD [jayasumana2024rethinkingfidbetterevaluation], and DreamSim [fu2023dreamsim], which compare the feature distributions or embeddings of generated and real images in pretrained perceptual spaces. Finally, we evaluate no-reference perceptual quality using MUSIQ [ke2021musiqmultiscaleimagequality] and NIQE [6353522], which estimate realism and naturalness without ground-truth supervision. While PSNR and SSIM directly measure pixel fidelity at the native resolution, the perceptual and distributional metrics internally downsample or resize images before feature extraction, reducing sensitivity to high-frequency details. As a result, improvements in fine-scale sharpness – central to super-resolution methods – may be underrepresented by these metrics, motivating our use of a broad set of complementary evaluations.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) enables high-quality novel view synthesis, motivating interest in generating higher-resolution renders than those available during training. A natural strategy is to apply super-resolution (SR) to low-resolution (LR) input views, but independently enhancing each image introduces multi-view inconsistencies, leading to blurry renders. Prior methods attempt to mitigate these inconsistencies through learned neural components, temporally consistent video priors, or joint optimization on LR and SR views, but all uniformly apply SR across every image. In contrast, our key insight is that close-up LR views may contain high-frequency information for regions also captured in more distant views and that we can use the camera pose relative to scene geometry to inform where to add SR content. Building on this insight, we propose SplatSuRe, a method that selectively applies SR content only in undersampled regions lacking high-frequency supervision, yielding sharper and more consistent results. Across Tanks & Temples, Deep Blending, and Mip-NeRF 360, our approach surpasses baselines in both fidelity and perceptual quality. Notably, our gains are most significant in localized foreground regions where higher detail is desired.
Figure 1: SplatSuRe trains a 3D Gaussian Splatting [14] model to produce sharp, high-resolution novel views from low-resolution inputs. By selectively leveraging high-frequency cues already present in low-resolution training views and applying super-resolution only where needed, our method delivers greater detail and multi-view consistency than prior approaches without any additional training.
@@ 1 Introduction
Novel view synthesis aims to render unseen viewpoints given a set of multi-view images. 3D Gaussian Splatting (3DGS) [14] enables real-time, photorealistic novel view synthesis by representing scenes as an explicit set of anisotropic Gaussians optimized through differentiable splatting. While 3DGS excels at efficiency and reconstruction quality, its performance is tightly coupled to the resolution of the training images. Models trained on low-resolution (LR) inputs lack access to high-frequency signals present in high-resolution (HR) views, resulting in blurry textures, over-smoothed surfaces, and aliasing artifacts when rendered at higher test-time resolutions [33].
When restricted to only LR views, a natural strategy is to apply super-resolution (SR) to enhance them before fitting a 3D model [6]. However, single-image SR operates independently on each view and frequently introduces view-dependent hallucinated textures [27]. When these inconsistent SR predictions are used as direct supervision, the resulting 3D optimization receives conflicting gradients across viewpoints, degrading model quality. Prior methods attempt to mitigate these inconsistencies through learned neural components, temporally consistent video priors, or joint optimization on LR and SR views [23, 25, 30, 10, 32]. However, these methods inject SR content uniformly across the image, regardless of whether a region actually benefits from generative detail or is already well-constrained by existing LR observations.
In contrast, the key observation motivating our work is that images of a scene do not sample 3D content uniformly. A low-resolution view captured up close often contains enough high-frequency detail to supervise rendering of more distant views that observe the same region only coarsely. This disparity in multi-view sampling implies that many views already receive high-frequency supervision from closer views, whereas others that do not have any closer views that provide sufficient higher-frequency information would benefit from SR guidance. Applying SR indiscriminately therefore introduces unnecessary inconsistencies in well-resolved regions.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [14] enables real-time novel view synthesis by representing scenes as sets of anisotropic Gaussians optimized through differentiable rasterization. While it achieves high rendering efficiency, models trained with low-resolution (LR) images suffer from aliasing artifacts when rendered at higher resolutions, since Gaussians optimized at coarse scales become undersampled. Mip-Splatting [33] mitigates this aliasing by applying scale-adaptive 3D and 2D filtering while preserving radiance energy across resolutions, similar in spirit to Mip-NeRF [1]. While achieving significant improvement over vanilla 3DGS, its rendering quality is still tied to the resolution of training views and blurring can occur at higher resolutions. In contrast, our method uses super-resolution to further inform high-resolution information.
@@
Super-resolution (SR) has long been applied to neural radiance fields [20] to enhance novel view synthesis quality [26, 11, 5, 24, 35, 10, 22]. More recently, SR has been extended to 3D Gaussian Splatting (3DGS). Several variants aim to recover high-resolution detail and improve multi-view consistency [6, 23, 25, 30] through residual feature learning [30], uncertainty modeling [25, 30], per-scene refinement [25] or video super-resolution [23]. Among these, SRGS [6] jointly optimizes Gaussian parameters using both LR ground-truth images and super-resolved views produced by a frozen single-image SR model. However, applying this enhancement uniformly across all regions does not eliminate the effect of inconsistencies introduced by super-resolution. S2Gaussian [25] focuses on sparse view reconstruction and proposes an inconsistency modeling module trained per-scene to reduce inconsistencies in SR images. SuperGaussian [23] applies a video SR network to frames rendered from a low-resolution 3DGS model, using the resulting temporally consistent sequence for retraining the 3D model. While these approaches improve fidelity, they either rely on additional neural components to enforce consistency or lack spatial adaptivity in the use of SR. In contrast, our method utilizes camera pose information relative to scene geometry to determine undersampled regions, and selectively applies super-resolution in those regions to enhance sharpness while maintaining fidelity.
@@
Diffusion methods offer complementary advances in 3D representation and enhancement tasks. Several works jointly optimize diffusion and 3D parameters for view-consistent generation and super-resolution [17, 36, 32, 2]. GaussianSR [32] distills information from 2D super-resolution models as a loss for training a 3DGS model. 3DSR [2] uses a 3DGS model to enforce 3D consistency in the diffusion process, fitting it multiple times through denoising diffusion steps. Diffusion priors have also been applied as post-processing to enhance renders from 3DGS models [18, 29, 28]. While these methods rely on large pretrained diffusion models to predict images that reduce 3D inconsistencies, we leverage explicit geometric relationships between cameras and scene structure to determine where generative detail is needed, which can then be added via any of these complementary methods.
@@
In addition to image and video-based models, SR can also be performed directly in 3D, circumventing multi-view inconsistencies. Geometric point-based networks [31, 19, 4, 3] locally upscale point clouds to recover fine-grained geometry. However, these approaches typically only upsample geometry, while texture resolution still relies on 2D methods. As before, our method is orthogonal to these techniques and can be combined with them to further improve reconstruction fidelity.
@@ 3 Background
3D Gaussian Splatting (3DGS) [14] models a scene as a set of 3D Gaussians, each represented by a mean position $\boldsymbol{\mu}\in\mathcal{R}^{3}$, per-axis scale $\boldsymbol{s}\in\mathcal{R}^{3}$, rotation vector $\boldsymbol{q}\in\mathcal{R}^{4}$, scalar opacity $o\in\mathcal{R}_{+}$, and view-dependent color $\boldsymbol{c}$, represented as a base color with spherical harmonic coefficients. Volumetric rendering with alpha blending is used to splat the Gaussians onto the image plane and the resulting color for pixel $\boldsymbol{p}$ is given by:
@@
where $N$ is the number of Gaussians intersecting the pixel ray, $\alpha^{i}(\boldsymbol{p})=o^{i}e^{-\frac{1}{2}(\boldsymbol{p}-\boldsymbol{\mu}^{i})^{T}(\boldsymbol{\Sigma}_{{2D}}^{i})^{-1}(\boldsymbol{p}-\boldsymbol{\mu}^{i})}$ is the contribution of the $i^{th}$ Gaussian, and $\boldsymbol{c}^{i}$ is its view-dependent color. The 2D covariance matrix $\boldsymbol{\Sigma}_{2D}$ is given by the EWA Splatting approximation [37]. Gaussians are sorted in depth order before splatting to ensure that transmittance is computed correctly.
@@ 3.2 Super-Resolution for Gaussian Splatting
When trained solely on low-resolution (LR) images, 3DGS models lack access to the high-frequency cues present in high-resolution (HR) views, leading to over-smoothed textures and incomplete fine structure. This limitation motivates integrating super-resolution (SR) models into the 3DGS training pipeline. SRGS [6] employs a frozen single-image SR model to generate SR views and jointly optimizes Gaussian parameters against both LR ground truth and SR outputs, supplying explicit high-frequency supervision that LR views alone cannot provide. However, SRGS applies SR uniformly across the entire image, even in regions that already receive reliable high-frequency supervision from nearby LR views. Since SR is applied independently to each image, the generated details are not necessarily multi-view consistent – injecting them everywhere can introduce geometric or texture inconsistencies, leading to averaging effects in the model that render as blurring.
@@ 4.1 Gaussian Fidelity Score
We use this ratio as an approximation for the Gaussian’s sampling frequency across views. A high ratio indicates that the Gaussian is sampled at varying frequencies, meaning some views observe it with high fidelity and can supervise the others. Conversely, a ratio close to one indicates uniform sampling, indicating that this region lacks any higher-frequency observations and requires SR to add generated details. This interpretation also holds for view regions where the Gaussian projects near its maximal radius, since no other views provide higher frequency information.
@@ 4.2 Super-Resolution Region Selection
$\mathcal{M}(t)$ $=\{\mathcal{G}^{i}\ |\ t=\operatorname*{argmax}_{t^{{}^{\prime}}\in T}r^{i}_{t^{{}^{\prime}}}\;\forall i\in\{1,\ldots,N\}\},$ (8)
@@
where Render()($\cdot$) denotes splatting the LR model by alpha-blending the specified per-Gaussian values, $\mathcal{G}$ is the set of all Gaussians, $\mathbf{score}_{\mathcal{G}}$ is the vector of Gaussian fidelity scores, and $\mathbf{1}_{\mathcal{M}(t)}(\mathcal{G})$ is an indicator that is 1 for Gaussians in $\mathcal{M}(t)$ and 0 otherwise. The first term in Equation 9 ensures that undersampled regions with low fidelity scores receive SR, while the second term ensures that areas observed most closely by the current view also receive SR because no alternative view provides higher resolution information. Finally, the weight map is normalized to ensure that the magnitude of the SR loss is consistent across views. As illustrated in Figure 4, this per-view weight map is high in regions requiring SR and low in regions already sufficiently supervised by LR views.
@@ 4.3 SplatSuRe Training Objective
In each training iteration, the model is rendered at the target high-resolution. The render $R_{HR}$ is downsampled to produce $R_{HR}\downarrow$ and compared with the LR ground truth $I_{LR}$, yielding the LR loss:
@@
Figure 5: Qualitative results on Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [33] and SRGS [6], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves finer details in text (red box on truck), high-frequency patterns (yellow box on carpet and green box on tray) and distant objects observed in other views (blue box on church mural). Notably, it reduces Gaussian artifacts (orange arrow) observed in other methods. Additional results in Appendix A.4.
@@ 5 Experiments
We integrate our SplatSuRe method into the 3D Gaussian Splatting (3DGS) codebase [14], adding modules for true Gaussian radius computation, weight map computation and rendering, and auxiliary loss terms. Experiments are conducted at $4\times$ super-resolution using StableSR [27] and using a ratio threshold $\tau{=}1.1$ to generate weight maps. Our choice of ratio threshold and SR model are ablated in Sections 7.1 and 7.2, respectively. We use $\lambda{=}0.2$ and $\gamma{=}0.4$ for our losses defined in Equations 10-12.
@@
Table 1: Quantitative results on Tanks & Temples [15]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. The best, second best and third best entries are highlighted. Our SplatSuRe method achieves the strongest results across most metrics.
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 0.669 19.41 0.350 71.58 2.013 0.0895 57.776 3.412
3DGS [14] + StableSR [27] 0.751 22.47 0.300 59.29 1.123 0.0667 56.748 4.945
Mip-Splatting [33] 0.767 23.10 0.303 52.46 1.137 0.0597 46.571 5.043
SRGS [6] + StableSR [27] 0.771 23.32 0.286 49.11 1.048 0.0535 55.209 4.633
Ours + StableSR [27] 0.784 23.81 0.272 37.72 1.040 0.0413 58.332 3.928
Table 2: Quantitative results on Deep Blending [8] and Mip-NeRF 360 [1]. Our SplatSuRe method achieves the strongest results across all metrics on Deep Blending and outperforms SRGS [6] on Mip-NeRF 360. Appendix A.1 and A.3 present $8\times$ SR and per-scene results.
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 0.836 26.72 0.335 0.868 0.0574 0.635 20.67 0.384 0.645 0.0529
3DGS [14] + StableSR [27] 0.845 27.16 0.325 0.630 0.0505 0.699 24.28 0.351 0.725 0.0387
Mip-Splatting [33] 0.865 28.43 0.327 0.690 0.0398 0.759 26.48 0.292 0.183 0.0132
SRGS [6] + StableSR [27] 0.861 28.23 0.317 0.630 0.0409 0.734 25.92 0.317 0.339 0.0194
Ours + StableSR [27] 0.872 29.01 0.306 0.496 0.0330 0.740 26.34 0.323 0.339 0.0179
@@
We compare SplatSuRe with four representative baselines using their official implementations on our evaluation datasets. 3DGS (LR) [14] is trained on LR inputs and is the primary baseline. 3DGS + StableSR [27] applies a single-image SR model to the low-resolution training data and fits a 3DGS model on the super-resolved images. Mip-Splatting [33] is trained at low-resolution and mitigates aliasing through 3D and 2D multi-scale filtering for high-resolution rendering. Finally, SRGS [6] optimizes Gaussian parameters using both low-resolution and super-resolved images from a frozen SR model, without additional pretraining or scene-specific fine-tuning.
@@
We evaluate both reconstruction fidelity and perceptual realism using eight complementary metrics. For reference-based assessment, we report SSIM, PSNR, and LPIPS [34], which measure pixel and feature-level agreement with ground-truth high-resolution images. To capture distributional and semantic perceptual quality, we include FID [9], CMMD [12], and DreamSim [7], which compare the feature distributions or embeddings of generated and real images in pretrained perceptual spaces. Finally, we evaluate no-reference perceptual quality using MUSIQ [13] and NIQE [21], which estimate realism and naturalness without ground-truth supervision. While PSNR and SSIM directly measure pixel fidelity at the native resolution, the perceptual and distributional metrics internally downsample or resize images before feature extraction, reducing sensitivity to high-frequency details. As a result, improvements in fine-scale sharpness – central to super-resolution methods – may be underrepresented by these metrics, motivating our use of a broad set of complementary evaluations.